OpenStudy (anonymous):

For the first 4 seconds, the displacement, in metres, of a sports car from its initial position is given by s=12t^2-t^3. The national speed limit in Great Britain is 70mph. Ath the end of 4 seconds, would the driver be breaking the law?

Parth (parthkohli):

lol, nice question by the way.

Parth (parthkohli):

The velocity of the car at time \(t\) is given by \(\dfrac{d}{dt}s\). Let's calculate what it is.

Parth (parthkohli):

\[v(t) = -3t^2 + 24t\]All righty, then. Now we have to calculate the maximum velocity the driver goes at. What is the max of \(v(t)\)?

OpenStudy (anonymous):

I dont understand this equation...\[t=ds/dt\] Could you please explain.

Parth (parthkohli):

Do you know that the derivative of displacement is velocity?

OpenStudy (anonymous):

no

Parth (parthkohli):

Do you know what derivative, displacement (or distance) and velocity (or speed) are?

OpenStudy (anonymous):

I know what they are, i just dont understand the 2 key relationships such as \[v=ds/dt \] and \[a=dv/dt\]

Parth (parthkohli):

All right. The derivative of something is just the rate-of-change of something. Do you understand that? If so, then it should be very easy to explain.

OpenStudy (anonymous):

yes

Parth (parthkohli):

And you know that, 1. Velocity is the rate of change of displacement. 2. Acceleration is the rate of change of velocity.

OpenStudy (anonymous):

yes

Parth (parthkohli):

Thus, we can conclude: 1. Velocity is the *derivative* of displacement. 2. Acceleration is the *derivative* of velocity. Got it? :D

OpenStudy (anonymous):

So , to work out the velocity, you differentiate the displacement

Parth (parthkohli):

Exactly. Similarly, for acceleration, you differentiate velocity.

OpenStudy (anonymous):

got it...

OpenStudy (anonymous):

So now could you answer the question plz...

OpenStudy (anonymous):

So what do you do after you differentiate the displacement getting \[-3t^2+24\]

Parth (parthkohli):

Great stuff! :) So \(v(t) = \dfrac{d}{dt}s(t) = -3t^2 + 24t\). You need to figure out whether \(-3t^2 + 24t\) goes above \(70\) in the interval \(0\le t \le 4\) because the displacement is given for the first four seconds.

OpenStudy (anonymous):

okay

Parth (parthkohli):

\(-3t^2 + 24t\) is the velocity, and the guy gets arrested if this function goes over 70. Can you figure out the maximum value of \(-3t^2 + 24\) in the interval \([0,4]\)?

OpenStudy (anonymous):

-24

OpenStudy (anonymous):

I mean 48

Parth (parthkohli):

Eek, I forgot one more thing: there is a difference between speed and velocity. Velocity is the absolute value of speed.

Parth (parthkohli):

Speed is the absolute value of velocity**

Parth (parthkohli):

Meaning that no matter what the velocity is, the speed will always be positive. If velocity = -24, then speed = 24.

Parth (parthkohli):

So we have to find the minimum of the function.

Parth (parthkohli):

If for example, the minimum velocity is -80, then the speed would be 80 which still exceeds the max speed.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

So what would the next step be...

OpenStudy (anonymous):

i Have worked out the answer. it is \[172.8 km h ^{-1}\], as i got the answer of \[48ms ^{-1}\] from differentiate the displacement and putting 4. Then i had to convert that number to miles per hour and so it was \[\frac{ 48 \times 60 \times 60 }{ 1000 }\], which equaled 172.8. So the person was faster than the national speed limit.